### v1

@@ 3.3 Magnetocentrifugal Winds
A number of authors in the 1970’s [134, 135] proposed that double-lobed radio sources could be the result of energy and angular momentum losses in a magnetohydrodynamic wind emerging from the upper and lower faces of an accretion disc. The dynamical structure of this wind was fully developed in the force-free limit by [134] and in magnetohydrodynamics by [136], with the disc as a lower boundary condition. Accretion via angular momentum losses in a magnetocentrifugal wind has received considerable attention in the context of weakly ionized flows, such as protoplanetary discs, where the interior of the disc may not be able to sustain MRI turbulence (see, e.g., [137] and references therein).
@@
Vertically stratified shearing box simulations with relatively strong vertical magnetic fields can in fact expel mass and angular momentum in an outflow from the upper and lower boundaries [138, 139, 118]. However, the outflow rates have a strong dependence on simulation box size [140], and global simulations are really essential to address this form of transport. Such global simulations, initialized with vertical magnetic fields, have been done [130, 141, 131, 132, 133], but the wind is generally subdominant in terms of outward angular momentum transport compared to magnetic stresses within the disc, at least in the simulations of [131, 132, 133]. It is worth noting that vertical magnetic fields are necessary but not sufficient for the formation of magnetocentrifugal winds. These winds are commonly observed in isothermal local shearing box simulations with vertical magnetic fields but not in radiation magneto-hydrodynamic local shearing box simulations with vertical magnetic fields where thermal properties of the gas are calculated self-consistently [142].
@@
An interesting feature of simulations with vertical magnetic flux and winds is that they exhibit a tendency to spontaneously form ring-like concentrations of surface density very much like the expected outcome of the Lightman-Eardley [2] viscous instability. Such “zonal flows”, where radial pressure gradients balance departures from Keplerian rotation (i.e. geostrophic balance between the Coriolis and pressure gradient forces in the local co-rotating frame) have also been seen in shearing box simulations without net vertical magnetic flux and without winds [143, 144]. However, the mechanism in that case appears to be associated with nonlinear self-organization in MRI turbulence [145]. In the presence of vertical magnetic fields, the fields end up concentrating in the low density “gaps” between the high mass density rings [146, 145, 133]. Possible explanations for this behavior are reconnection within vertical MRI channel modes that cause separation of mass from vertical flux [146] or a true linear viscous instability caused by the winds [145]. Neither of these proposals appear to explain the simulation results of [133], who also note that the rings disappear above some critical threshold of vertical magnetization.
@@
The most compelling application of thermal/viscous instabilities in accretion discs to observational reality is in outbursting compact binaries. There exist quite complete reviews describing the disc instability model of these outbursts, one recent [147] and two more ancient (Osaki 148, Lasota 149), so in the present section we will present and discuss the basics of the model, and supplemented only by the newest results and open questions.
@@ 4.1 Dwarf novae
Cataclysmic variable stars (CVs; see Warner 150) can be divided into two types: one consists of quasi-steady systems, known as “nova-likes” (NLs), and the other consists of systems that are strongly variable on various timescales and are known as dwarf novae (DNs). It is now well established that the two sets of binaries differ by the nature of their accretion discs. The membership of one of the two classes of CVs is determined by two physical parameters: the mass-transfer rate $\dot{M}_{\rm tr}$ from the low-mass stellar companion of the white dwarf (WD) and the size $R_{\rm D}$ of the accretion disc around the WD. These parameters correspond to observed quantities: luminosity and orbital period. The physical reason for the separation of CVs into two classes is a thermal instability affecting accretion discs at temperatures below $\sim 10^{4}$ K. Such discs also suffer from a viscous instability but since in geometrically thin discs the thermal time is much shorter than the viscous time, the disc variability is triggered by thermal processes.
Stability of the discs has been studied by small perturbation methods. Linearized perturbations of the mass transfer and energy balance equations for a geometrically thin accretion disc by Piran [4] provided two necessary conditions for a stable disc. Consider an $\alpha$-disc, with a viscous heating rate $Q^{+}$ and a radiative cooling rate $Q^{-}=\sigma\,T_{\rm eff}^{4}$. The thermal and viscous stability conditions are, respectively:
@@
The thermal-viscous instability of $\alpha$-discs, when both conditions (22) are violated, is associated with hydrogen partial ionization. It is a result of (i) the opacity in the outer layers of the disc at temperatures $\lesssim 10^{4}$ K dramatically drops and strongly depends on the temperature, $d\ln\kappa_{\rm R}/d\ln T\approx 7-10$; and (ii) the surface boundary condition dominates the structure [151]. In particular, this is why irradiation can suppress instability by maintaining the temperature only at the disc surface.
Figure 2: Mass transfer rates of CVs compared to the stability criterion Eq. (25). Systems above the upper (red) solid line are hot and stable. Systems below the lower (blue) line indicate cold, stable discs if the white dwarf magnetic field $B\geq 10^{5}$ G. The dashed line represents the expected secular mass transfer rate (Knigge et al. 152). Square symbols indicate Z Cam type dwarf novæ; (red) stars indicate nova-likes. Dwarf novae shown in black have a more complete observed light-curve than those in grey.
@@
for every value of the accretion rate there exists a size of a disc at which it becomes unstable. One could get the $\dot{M}_{\rm crit}(R_{\rm D})$ critical relation from Eq. (23) by putting there $T_{\rm eff}=6500$ K (see e.g. Smak 153), the temperature corresponding to hydrogen recombination, but more detailed calculations give the relation [154]
@@
The methods used to obtain reliable $\dot{M}$ from observed CV magnitudes are described in detail in Dubus et al. [155]. The disc radius is expressed through the orbital period by assuming that it is a fraction $f(q)$ of the binary separation $a$: $R_{\rm D}=f(q)a=3.5\times 10^{10}f(q)M_{1}^{1/3}P^{2/3}_{\rm h}\rm cm$, where $q$ is the mass ratio (mass-of-the-companion/white-dwarf-mass) and $P_{\rm h}$ the orbital period in hours. In general, $f$ is well approximated by $f=0.6/(1+q)^{2/3}$. The critical accretion rate can then be written as
@@
Figure 2 [155] exhibiting the $\dot{M}\left(P_{\rm orb}\right)$ relation deduced from observations of about 130 CVs shows clearly that the relation Eq. (26) separates these binary systems into two classes: those above are NLs while those below are DNs.
@@
Figure 3: $\Sigma-T_{\rm eff}$ S-curves computed for a $1.35M_{\odot}$ white dwarf are plotted for various radii. The four sets of curves (each set includes a dashed, a solid and a dotted line), from left to right, represent the S-curves at radii $R=10^{9},10^{10},10^{11}$ and $10^{12}\rm cm$, respectively. Solid lines represent the S-curves obtained with temperature-dependent $\alpha$. Dashed and dotted lines represent S-curves for $\alpha=0.1$ and $\alpha=0.01$ respectively [156]).
@@ 4.2 The dwarf-nova disc instability model
where $\nu$ is the kinematic viscosity coefficient. Since $\nu=\alpha\,c_{s}^{2}\,\Omega_{\scriptscriptstyle{\mathrm{K}}}^{-1}\propto T_{\rm c}$, where $c_{s}$ is the isothermal speed of sound, solving the vertical energy transfer equation provides a relation between the effective and midplane temperatures. This allows one to represent the disc equilibria as a $T_{\rm eff}(\Sigma)$ or $Q(\Sigma)$ relation. As discovered in the early 1980’s this relation forms an S – a necessary condition for the presence of outbursts forming a limit cycle [157]. The upper, hot branch of the S corresponds to disc states in outburst while during quiescence they are located on the lower, cold part of the curve. The middle part of the S is thermally and viscously unstable. This is directly related to the fact that the slope of the curve $Q(\Sigma)$ equals the positive or negative number $(k-l)/(k-1)$, for the form of $Q^{-}$ suggested by (22).
The states on both stable branches evolve on a viscous time and cross, up and down, the space between the two in a thermal time. Since, in general, the DN outburst duration is shorter than the quiescence time, if the limit cycle is to reproduce the DN eruption cycle, the viscous time on the upper branch of the S-curve should be shorter than the viscous time on the cold segment of the S. Since the viscous time $t_{\rm vis}\sim T_{c}^{-1}$ this indeed is the case but as it has been quickly discovered, the difference due to the temperature alone is not sufficient to produce a light curve resembling a usual DN outburst cycle [158].
The viscosity parameter $\alpha$ for hot discs has been determined from observations of DN outburst decays to be $\alpha_{h}=0.1-0.2$ [159, 160]. To reproduce the observed outbursts amplitudes, $\alpha_{c}$ in cold discs must be 5 to 10 times smaller. These two conclusions are rather embarrassing from the theoretical point view because, until recently, MRI simulations with no net vertical flux usually produced $\alpha_{h}$ ten times smaller and the value of $\alpha_{c}$ is difficult to estimate. These problems will be discussed in more detail in §4.3. Here we should stress that despite its shaky theoretical background the “two - $\alpha$’s paradigm” is very successful in reproducing many properties of the DN outbursts.
In practice this paradigm consists of joining, at a given radius, two S-curves into one “effective” S-curve with different values of $\alpha$ on the hot and cold branches (see e.g. Hameury and Lasota 161 Eq. 17). The results of such an operation are shown in Fig. 3.
Figure 4: Left: the lightcurve of SS Cyg observed by the AAVSO; Right: three cases of model lightcurves [162] of a binary with SS Cyg parameters. From the top: for $R_{\rm in}=R_{\rm WD}$ and $\dot{M}_{\rm tr}=const.$; $R_{\rm in}>R_{\rm WD}$ and $\dot{M}_{\rm tr}=const.$, and $R_{\rm in}>R_{\rm WD}$ and $\dot{M}_{\rm tr}\neq const.$
@@ 4.3 MHD Models of the Ionization Instability in Compact Binaries
As we just discussed in the context of alpha-viscosities in the disc instability model, the outbursting behavior in cataclysmic variables provides the best observational constraints on the physical mechanisms of angular momentum transport in accretion discs. Two $\alpha$’s are needed: a high value $\sim 0.1-0.2$ in the hot outburst state, and a much lower value in the cold quiescent state. For hydrogen accreting systems, both hydrogen and helium are largely neutral in the quiescent state, and the dominant sources of free electrons are from alkali metals (particularly sodium) which have first ionization potentials less than that of hydrogen. The abundance of these elements is rather low, however, implying that the plasma in the quiescent state would be too resistive to support robust MRI turbulence. This alone might explain why $\alpha_{c}$ is so low in the quiescent state of hydrogen accreting systems, and might point to an alternative mechanism whereby it is the onset of good conductivity and MRI turbulence, not thermal instability per se, that drives outbursts [163]. Even if that were the case, however, there remains the problem of the high values of $\alpha_{h}$ in the outburst state, which are an order of magnitude larger than the values inferred from local shearing box simulations of MRI turbulence with no net vertical flux (§3.1). It may be that the discrepancy arises because of the limitations of the shearing box, and that more global magnetic field structures that can link different radii together are necessary to enhance the angular momentum transport [164, 165].
@@
Despite these issues, progress has been made by incorporating the thermodynamics associated with changes in opacity and cooling rates in shearing box simulations of MRI turbulence. The first attempt at this [166] used vertically unstratified shearing boxes with an optically thin cooling function which mocked up the expected cooling rates across the ionization transition. Thermally stable equilibria at high and low temperatures were found, with unstable behavior in between, replicating the bistable behavior found in the disc instability model. Vertically stratified shearing boxes with diffusive radiation transport and realistic grey opacities have also been done [112, 115], and found thermally stable equilibria at low and high temperatures (as well as a stable intermediate temperature branch in the case of [115]), with unstable behavior in between. An example is shown in Figure 5. In other words, simulations with MRI turbulence and realistic cooling are able to replicate stable branches of the S-curves used in the disc instability model. Moreover, the vertically stratified simulations also find high values of alpha in the hot stable branch, comparable to those inferred from observations, even without net vertical flux in these simulations. The major difference with the standard disc instability model is that the enhanced values of alpha are only found near the low temperature end of the upper branch: the alpha’s return to lower values characteristic of standard zero net vertical flux simulations higher up on the upper branch. The reason for the enhanced alpha’s at the end of the branch are because the high opacities drive thermal convection, and this enhances the time-averaged MRI stresses (see §3.1). Some of the simulations by [115] incorporated resistivity and confirmed the conclusion of [163] that MRI turbulence would be suppressed over much, but not all, of the lower branch - the high surface density end of the branch can still be sufficiently conducting to sustain MRI turbulence.
@@
Figure 5: Stable thermal equilibria in vertically stratified shearing box simulations of MRI turbulence with diffusive radiation transport (from [112]) and realistic opacities for a solar-composition plasma. The colors indicate the time-averaged values of the Shakura-Sunyaev $\alpha$-parameter as measured from the MRI turbulent stresses and thermal pressures. Grey curves show the thermal equilibrium predictions of the standard disc instability model with constant $\alpha$.
@@
Such MRI-based simulations of S-curves have also been done in the context of protoplanetary discs [113] (with possible implications for FU Ori outbursts) and helium-dominated AM CVn discs [167], which exhibit both normal and superoutbursts. Once again, convection enhances the effective $\alpha$ parameter at the lower end of the upper stable branch. One interesting implication of the helium-dominated simulations is that convection can be persistent and still lead to enhanced $\alpha$ parameter near the end of the upper branch. Moreover, because the first ionization potential of helium is so high, there are much more abundant sources of free electrons (particularly nitrogen) in the cool branch which lead to high enough electrical conductivity that ideal MHD should be a good approximation [167]. It would therefore be interesting to explore the differences between AM CVn outbursts and hydrogen-dominated dwarf novae as they might relate to these differences in conductivity in the quiescent state.
@@
Whether merely enhancing the value of $\alpha$ at the low end of the upper branch is sufficient to explain outburst light curves is not as yet clear. Ideally, one would like to run global radiation MHD simulations of MRI turbulence with realistic opacities to follow the evolution of heating and cooling fronts through an outburst cycle, but this remains numerically challenging. An attempt was made to incorporate the variation of $\alpha$ on the upper branch into a dwarf nova outburst code [168] to compute outburst light curves, and these did replicate observed outburst and quiescent state time scales, as well as outburst amplitudes. However, they also produced reflares in the decay to quiescence that are not generally observed. A contributing factor to this was the poor modeling of the quiescent state [168], where as we just discussed may in any case be too resistive to sustain MRI turbulence. Much more work needs to be done to explore the physics of the quiescent state.
@@
Radiation MHD simulations in vertically stratified shearing boxes with net vertical magnetic flux have also been pursued [169]. Convection can again enhance $\alpha$ in the upper branch, but when the vertical magnetic field is strong enough, it can itself cause an increase in $\alpha$, and this enhancement can also occur in the quiescent state. Vertical magnetic flux can also extract angular momentum in a magnetocentrifugal wind. Incorporating such wind-driven transport into the disc instability model can replicate observed outburst lightcurves [170, 171].
@@
Before leaving this topic, it is also worth mentioning that, while global radiation MHD simulations of outbursts are not yet practicable, such simulations with MRI turbulence and no radiation transport have been pursued in Roche potentials. These are shedding light on the relative roles of spiral waves and MRI turbulence in angular momentum transport [77, 78, 172], and how discs may spread outward in such potentials [173, 174]. They have also recently been able to replicate the growth of eccentricity in the presence of MRI turbulence due to the 3:1 mean motion resonance [174], which is thought to be the origin of observed (positive) superhumps [175, 176].

### v2

@@ 3.3 Magnetocentrifugal Winds
A number of authors in the 1970’s [bla76, lov76] proposed that double-lobed radio sources could be the result of energy and angular momentum losses in a magnetohydrodynamic wind emerging from the upper and lower faces of an accretion disc. The dynamical structure of this wind was fully developed in the force-free limit by bla76 and in magnetohydrodynamics by bla82, with the disc as a lower boundary condition. Accretion via angular momentum losses in a magnetocentrifugal wind has received considerable attention in the context of weakly ionized flows, such as protoplanetary discs, where the interior of the disc may not be able to sustain MRI turbulence [see, e.g., les21, and references therein].
@@
Vertically stratified shearing box simulations with relatively strong vertical magnetic fields can in fact expel mass and angular momentum in an outflow from the upper and lower boundaries [suz09, les13, bai13]. However, the outflow rates have a strong dependence on simulation box size [fro13], and global simulations are really essential to address this form of transport. Such global simulations, initialized with vertical magnetic fields, have been done [suz14, ava16, zhu18, mis20, jac21], but the wind is generally subdominant in terms of outward angular momentum transport compared to magnetic stresses within the disc, at least in the simulations of zhu18, mis20, jac21. It is worth noting that vertical magnetic fields are necessary but not sufficient for the formation of magnetocentrifugal winds. These winds are commonly observed in isothermal local shearing box simulations with vertical magnetic fields but not in radiation magneto-hydrodynamic local shearing box simulations with vertical magnetic fields where thermal properties of the gas are calculated self-consistently [Secunda+2024].
@@
An interesting feature of simulations with vertical magnetic flux and winds is that they exhibit a tendency to spontaneously form ring-like concentrations of surface density very much like the expected outcome of the lig74 viscous instability. Such “zonal flows”, where radial pressure gradients balance departures from Keplerian rotation (i.e. geostrophic balance between the Coriolis and pressure gradient forces in the local co-rotating frame) have also been seen in shearing box simulations without net vertical magnetic flux and without winds [joh09, sim12]. However, the mechanism in that case appears to be associated with nonlinear self-organization in MRI turbulence [rio19]. In the presence of vertical magnetic fields, the fields end up concentrating in the low density “gaps” between the high mass density rings [bai14, rio19, jac21]. Possible explanations for this behavior are reconnection within vertical MRI channel modes that cause separation of mass from vertical flux [bai14] or a true linear viscous instability caused by the winds [rio19]. Neither of these proposals appear to explain the simulation results of jac21, who also note that the rings disappear above some critical threshold of vertical magnetization.
@@
The most compelling application of thermal/viscous instabilities in accretion discs to observational reality is in outbursting compact binaries. There exist quite complete reviews describing the disc instability model of these outbursts, one recent [Hameury0420] and two more ancient (Osaki01996, Lasota0601), so in the present section we will present and discuss the basics of the model, and supplemented only by the newest results and open questions.
@@ 4.1 Dwarf novae
Cataclysmic variable stars (CVs; see Warner2003) can be divided into two types: one consists of quasi-steady systems, known as “nova-likes” (NLs), and the other consists of systems that are strongly variable on various timescales and are known as dwarf novae (DNs). It is now well established that the two sets of binaries differ by the nature of their accretion discs. The membership of one of the two classes of CVs is determined by two physical parameters: the mass-transfer rate $\dot{M}_{\rm tr}$ from the low-mass stellar companion of the white dwarf (WD) and the size $R_{\rm D}$ of the accretion disc around the WD. These parameters correspond to observed quantities: luminosity and orbital period. The physical reason for the separation of CVs into two classes is a thermal instability affecting accretion discs at temperatures below $\sim 10^{4}$ K. Such discs also suffer from a viscous instability but since in geometrically thin discs the thermal time is much shorter than the viscous time, the disc variability is triggered by thermal processes.
Linearized perturbations of the mass transfer and energy balance equations for a geometrically thin accretion disc by pir78 provided two necessary conditions for a stable disc. Consider an $\alpha$-disc, with a viscous heating rate $Q^{+}$ and a radiative cooling rate $Q^{-}=\sigma\,T_{\rm eff}^{4}$. The thermal and viscous stability conditions are, respectively:
@@
The thermal-viscous instability of $\alpha$-discs, when both conditions (22) are violated, is associated with hydrogen partial ionization. It is a result of (i) the opacity in the outer layers of the disc at temperatures $\lesssim 10^{4}$ K dramatically drops and strongly depends on the temperature, $d\ln\kappa_{\rm R}/d\ln T\approx 7-10$; and (ii) the surface boundary condition dominates the structure [Faulkner_etal1983_1]. In particular, this is why irradiation can suppress instability by maintaining the temperature only at the disc surface.
Figure 5: Mass transfer rates of CVs compared to the stability criterion Eq. (25). Systems above the upper (red) solid line are hot and stable. Systems below the lower (blue) line indicate cold, stable discs if the white dwarf magnetic field $B\geq 10^{5}$ G. The dashed line represents the expected secular mass transfer rate (Knigge0611). Square symbols indicate Z Cam type dwarf novæ; (red) stars indicate nova-likes. Dwarf novae shown in black have a more complete observed light-curve than those in grey.
@@
for every value of the accretion rate there exists a size of a disc at which it becomes unstable. One could get the $\dot{M}_{\rm crit}(R_{\rm D})$ critical relation from Eq. (23) by putting there $T_{\rm eff}=6500$ K (see, e.g., Smak0182), the temperature corresponding to hydrogen recombination, but more detailed calculations give the relation [Lasota0808]
@@
The methods used to obtain reliable $\dot{M}$ from observed CV magnitudes are described in detail in Dubus0918. The disc radius is expressed through the orbital period by assuming that it is a fraction $f(q)$ of the binary separation $a$: $R_{\rm D}=f(q)a=3.5\times 10^{10}f(q)M_{1}^{1/3}P^{2/3}_{\rm h}\rm cm$, where $q$ is the mass ratio (mass-of-the-companion/white-dwarf-mass) and $P_{\rm h}$ the orbital period in hours. In general, $f$ is well approximated by $f=0.6/(1+q)^{2/3}$. The critical accretion rate can then be written as
@@
Fig. 5 [Dubus0918] exhibiting the $\dot{M}\left(P_{\rm orb}\right)$ relation deduced from observations of about 130 CVs shows clearly that the relation Eq. (26) separates these binary systems into two classes: those above are NLs while those below are DNs.
@@
Figure 6: $\Sigma-T_{\rm eff}$ S-curves computed for a $1.35M_{\odot}$ white dwarf are plotted for various radii. The four sets of curves (each set includes a dashed, a solid and a dotted line), from left to right, represent the S-curves at radii $R=10^{9},10^{10},10^{11}$ and $10^{12}\rm cm$, respectively. Solid lines represent the S-curves obtained with temperature-dependent $\alpha$. Dashed and dotted lines represent S-curves for $\alpha=0.1$ and $\alpha=0.01$ respectively [Bollimpalli1218].
@@ 4.2 The dwarf-nova disc instability model
where $\nu$ is the kinematic viscosity coefficient. Since $\nu=\alpha\,c_{s}^{2}\,\Omega_{\scriptscriptstyle{\mathrm{K}}}^{-1}\propto T_{\rm c}$, where $c_{s}$ is the isothermal speed of sound, solving the vertical energy transfer equation provides a relation between the effective and midplane temperatures. This allows one to represent the disc equilibria as a $T_{\rm eff}(\Sigma)$ or $Q(\Sigma)$ relation. As discovered in the early 1980’s this relation forms an S – a necessary condition for the presence of outbursts forming a limit cycle [Bath0482]. The upper, hot branch of the S corresponds to disc states in outburst while during quiescence they are located on the lower, cold part of the curve. The middle part of the S is thermally and viscously unstable. This is directly related to the fact that the slope of the curve $Q(\Sigma)$ equals the positive or negative number (kl)/(k1)(k-l)/(k-1), for the form of $Q^{-}$ suggested by (22).
The states on both stable branches evolve on a viscous time and cross, up and down, the space between the two in a thermal time. Since, in general, the DN outburst duration is shorter than the quiescence time, if the limit cycle is to reproduce the DN eruption cycle, the viscous time on the upper branch of the S-curve should be shorter than the viscous time on the cold segment of the S. Since the viscous time $t_{\rm vis}\sim T_{c}^{-1}$ this indeed is the case but as it has been quickly discovered, the difference due to the temperature alone is not sufficient to produce a light curve resembling a usual DN outburst cycle [Smak0184].
The viscosity parameter $\alpha$ for hot discs has been determined from observations of DN outburst decays to be $\alpha_{h}=0.1-0.2$ [Smak0999, Kotko0912]. To reproduce the observed outbursts amplitudes, $\alpha_{c}$ in cold discs must be 5 to 10 times smaller. These two conclusions are rather embarrassing from the theoretical point view because, until recently, MRI simulations with no net vertical flux usually produced $\alpha_{h}$ ten times smaller and the value of $\alpha_{c}$ is difficult to estimate. These problems will be discussed in more detail in §4.3. Here we should stress that despite its shaky theoretical background the “two - $\alpha$’s paradigm” is very successful in reproducing many properties of the DN outbursts.
In practice this paradigm consists of joining, at a given radius, two S-curves into one “effective” S-curve with different values of $\alpha$ on the hot and cold branches [see, e.g., Hameury0485, Eq. 17]. The results of such an operation are shown in Fig. 6.
Figure 7: Left: the lightcurve of $S$ Cyg observed by the AAVSO; Right: three cases of model lightcurves [Schreiber1003] of a binary with $S$ Cyg parameters. From the top: for $R_{\rm in}=R_{\rm WD}$ and $\dot{M}_{\rm tr}=const.$; $R_{\rm in}>R_{\rm WD}$ and $\dot{M}_{\rm tr}=const.$, and $R_{\rm in}>R_{\rm WD}$ and $\dot{M}_{\rm tr}\neq const.$
@@ 4.3 MHD Models of the Ionization Instability in Compact Binaries
As we just discussed in the context of alpha-viscosities in the disc instability model, the outbursting behavior in cataclysmic variables provides the best observational constraints on the physical mechanisms of angular momentum transport in accretion discs. Two $\alpha$’s are needed: a high value $\sim 0.1-0.2$ in the hot outburst state, and a much lower value in the cold quiescent state. For hydrogen accreting systems, both hydrogen and helium are largely neutral in the quiescent state, and the dominant sources of free electrons are from alkali metals (particularly sodium) which have first ionization potentials less than that of hydrogen. The abundance of these elements is rather low, however, implying that the plasma in the quiescent state would be too resistive to support robust MRI turbulence. This alone might explain why $\alpha_{c}$ is so low in the quiescent state of hydrogen accreting systems, and might point to an alternative mechanism whereby it is the onset of good conductivity and MRI turbulence, not thermal instability per se, that drives outbursts [gam98]. Even if that were the case, however, there remains the problem of the high values of $\alpha_{h}$ in the outburst state, which are an order of magnitude larger than the values inferred from local shearing box simulations of MRI turbulence with no net vertical flux (§3.1). It may be that the discrepancy arises because of the limitations of the shearing box, and that more global magnetic field structures that can link different radii together are necessary to enhance the angular momentum transport [tou96, kin07].
@@
Despite these issues, progress has been made by incorporating the thermodynamics associated with changes in opacity and cooling rates in shearing box simulations of MRI turbulence. The first attempt at this [lat12] used vertically unstratified shearing boxes with an optically thin cooling function which mocked up the expected cooling rates across the ionization transition. Thermally stable equilibria at high and low temperatures were found, with unstable behavior in between, replicating the bistable behavior found in the disc instability model. Vertically stratified shearing boxes with diffusive radiation transport and realistic grey opacities have also been done [hir14, sce18a], and found thermally stable equilibria at low and high temperatures [as well as a stable intermediate temperature branch in the case of sce18a], with unstable behavior in between. An example is shown in Fig. 8. In other words, simulations with MRI turbulence and realistic cooling are able to replicate stable branches of the S-curves used in the disc instability model. Moreover, the vertically stratified simulations also find high values of alpha in the hot stable branch, comparable to those inferred from observations, even without net vertical flux in these simulations. The major difference with the standard disc instability model is that the enhanced values of alpha are only found near the low temperature end of the upper branch: the alpha’s return to lower values characteristic of standard zero net vertical flux simulations higher up on the upper branch. The reason for the enhanced alpha’s at the end of the branch are because the high opacities drive thermal convection, and this enhances the time-averaged MRI stresses (see §3.1). Some of the simulations by sce18a incorporated resistivity and confirmed the conclusion of gam98 that MRI turbulence would be suppressed over much, but not all, of the lower branch — the high surface density end of the branch can still be sufficiently conducting to sustain MRI turbulence.
@@
Figure 8: Stable thermal equilibria in vertically stratified shearing box simulations of MRI turbulence with diffusive radiation transport from [from hir14] and realistic opacities for a solar-composition plasma. The colors indicate the time-averaged values of the Shakura-Sunyaev $\alpha$-parameter as measured from the MRI turbulent stresses and thermal pressures. Grey curves show the thermal equilibrium predictions of the standard disc instability model with constant $\alpha$.
@@
Such MRI-based simulations of S-curves have also been done in the context of protoplanetary discs [hir15, with possible implications for FU Ori outbursts] and helium-dominated AM CVn discs [col18], which exhibit both normal and superoutbursts. Once again, convection enhances the effective $\alpha$ parameter at the lower end of the upper stable branch. One interesting implication of the helium-dominated simulations is that convection can be persistent and still lead to enhanced $\alpha$ parameter near the end of the upper branch. Moreover, because the first ionization potential of helium is so high, there are much more abundant sources of free electrons (particularly nitrogen) in the cool branch which lead to high enough electrical conductivity that ideal MHD should be a good approximation [col18]. It would therefore be interesting to explore the differences between AM CVn outbursts and hydrogen-dominated dwarf novae as they might relate to these differences in conductivity in the quiescent state.
@@
Whether merely enhancing the value of $\alpha$ at the low end of the upper branch is sufficient to explain outburst light curves is not as yet clear. Ideally, one would like to run global radiation MHD simulations of MRI turbulence with realistic opacities to follow the evolution of heating and cooling fronts through an outburst cycle, but this remains numerically challenging. An attempt was made to incorporate the variation of $\alpha$ on the upper branch into a dwarf nova outburst code [col16] to compute outburst light curves, and these did replicate observed outburst and quiescent state time scales, as well as outburst amplitudes. However, they also produced reflares in the decay to quiescence that are not generally observed. A contributing factor to this was the poor modeling of the quiescent state [col16], where as we just discussed may in any case be too resistive to sustain MRI turbulence. Much more work needs to be done to explore the physics of the quiescent state.
@@
Radiation MHD simulations in vertically stratified shearing boxes with net vertical magnetic flux have also been pursued [sce18b]. Convection can again enhance $\alpha$ in the upper branch, but when the vertical magnetic field is strong enough, it can itself cause an increase in $\alpha$, and this enhancement can also occur in the quiescent state. Vertical magnetic flux can also extract angular momentum in a magnetocentrifugal wind. Incorporating such wind-driven transport into the disc instability model can replicate observed outburst lightcurves [sce19, sce20].
@@
Before leaving this topic, it is also worth mentioning that, while global radiation MHD simulations of outbursts are not yet practicable, such simulations with MRI turbulence and no radiation transport have been pursued in Roche potentials. These are shedding light on the relative roles of spiral waves and MRI turbulence in angular momentum transport [ju16, ju17, pja20], and how discs may spread outward in such potentials [oya21, oha24]. They have also recently been able to replicate the growth of eccentricity in the presence of MRI turbulence due to the 3:1 mean motion resonance [oha24], which is thought to be the origin of observed (positive) superhumps [whi88, lub91].